I would really appreciate your help, I am very stumped on this one!!! It says to multiply and simplify by factoring!!
radical equations 4√16 * 4√32 Both of the 4's are up in the checkmark. The fourth root I believe it what it is called..

Answer by stanbon(75887)

It says to multiply and simplify by factoring!!
radical equations 4√16 * 4√32 Both of the 4's are up in the checkmark. The fourth root I believe it what it is called..
----------------------
16 factors as 2^4
32 factors as 2^5
--------------------------
4th rt. of 16 = 4th rt. of 2^4 = 2
---
4th rt. of 32 = 4th rt. of 2^5 = (4th rt of 2^4) times 2 = 2*(4th rt of 2)
----------------
So you get 2*2*4th rt of 2 = 4*4th rt of 2 as an answer.
==============
Cheers,
Stan H.